Solve the problem If f(x) = x2 - 2 and g(x) = x + 3, find (f o g)(4) and (go (-4). (fog)(4) = (go(-4)
Answer #1

fo= 82 9cz (443) (fogo Os z f(x+3) z (x+33-2 fog & 2 (4+3,²_2 = 72_2 fogjes = 49-2-47 | stog) ch=47 (gof) 2012 g(+6)= g(8-2)